OpenStudy (anonymous):

(05.09) Reggie can line a football field in 120 minutes. Rosalinda can line a football field in 80 minutes. If they work together, how many minutes does it take them to line a football field?

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

\[\frac{ x }{ 120 } + \frac{ x }{ 80 } = 1\]

OpenStudy (anonymous):

you can use this formula: \[\frac{ 1 }{ t }=\frac{ 1 }{ a }+\frac{ 1 }{ b }\] in this formula a is the time that first person can do a work. b is the time that second person can do a work. t is the time that if both persons work together can finish it. \[\frac{ 1 }{ t }=\frac{ 1 }{ 80 }+\frac{ 1 }{ 120 }\]